# A Deep Probabilistic Spatiotemporal Framework for Dynamic Graph Representation Learning with Application to Brain Disorder Identification

### v3

@@ Abstract
Recent applications of pattern recognition techniques on brain connectome classification using functional connectivity (FC) are shifting towards acknowledging the non-Euclidean topology and causal dynamics of brain connectivity across time. In this paper, a deep spatiotemporal variational Bayes (DSVB) framework is proposed to learn time-varying topological structures in dynamic FC networks for identifying autism spectrum disorder (ASD) in human participants. The framework incorporates a spatial-aware recurrent neural network with an attention-based message passing scheme to capture rich spatiotemporal patterns across dynamic FC networks. To overcome model overfitting on limited training datasets, an adversarial training strategy is introduced to learn graph embedding models that generalize well to unseen brain networks. Evaluation on the ABIDE resting-state functional magnetic resonance imaging dataset shows that our proposed framework substantially outperforms state-of-the-art methods in identifying patients with ASD. Dynamic FC analyses with DSVB-learned embeddings reveal apparent group differences between ASD and healthy controls in brain network connectivity patterns and switching dynamics of brain states.
@@ 1 Introduction
In this paper, a deep spatiotemporal variational Bayes (DSVB) framework is proposed to learn the stochastic dynamic graph embeddings of FC networks, composed of sliding window segments of the resting-state fMRI time-series data. Figure 1 shows an overview of the proposed DSVB framework. Derived from general sequential variational Bayes (VB), our proposed framework provides a stochastic representation modelling solution of using dynamic FC networks for neuropsychiatric disorder diagnosis. Unlike existing GNN-based approaches that learn independent and identically distributed (i.i.d.) graph embeddings, we incorporate a spatial-aware recurrent neural network with attention-based message-passing scheme to accurately model both the complex topological structure and causal dependencies within a dynamic FC network sequence. Additionally, we introduce an adversarial training strategy to overcome the overfitting issue aggravated by limited training data. Our contributions are highlighted as follows:
@@
    An end-to-end probabilistic framework, integrating variational Bayes and recurrent GNNs, to learn the sequential stochastic graph latent embedding spaces in a purely unsupervised manner. This accommodates a wider range of spatiotemporal variability across unseen FC networks.
@@ 2.2 Spatiotemporal Variational Bayes
where $\mathbf{X}_{\leq t}$ and $\mathbf{X}_{<t}$ denote the partial sequences up to the $t^{\text{th}}$ and $({t-1})^{\text{th}}$ time samples, respectively. $\mathcal{D}^{\mathrm{KL}}$ denotes the (positive-valued) Kullback–Leibler divergence (KLD), which measures the statistical discrepancy of the encoder from the prior. The conditional probabilities of (3) encapsulate the underlying causal structure and temporal coherence of the dynamic brain networks. This sELBO (3) then forms the basis of our proposed DSVB framework.
@@ 2.5 Latent Embeddings for Graph Classification
modelled by multilayer FCNN. A softmax activation is then applied on the obtained logits $\mathbf{\hat{y}}\in\mathbb{R}^{2}$ to obtain the predictive probability scores for the final subject-level brain network classification. The predicted class label is thus the one with highest predictive score. The classification loss is taken to be the following multi-class cross entropy:
@@ 4 Results and Discussion
To evaluate the effectiveness of the proposed method, a nested-stratified 5-fold cross-validation was applied for each experiment. The performance of ASD identification is measured using five metrics: accuracy, recall, precision, F1-score, and area under the curve (AUC). Table 1 compares the average performances of the proposed DSVB-FCNN with the existing baseline classifiers based on both dynamic and static FC networks. It shows that our proposed DSVB-FCNN achieves substantial improvement over the baseline classifiers in each metric. The DSVB-FCNN also achieved low standard deviations in all metrics, demonstrating a strong consistency in its performances across the 5-fold cross-validation. In general, these results indicate that the framework facilitates robust brain disorder identification by generating graph latent embeddings that generalize well to unseen dynamic FC networks beyond the limited training dataset.
@@
Figure 3: The average connectivity patterns derived from DSVB-learned embeddings. (a) 3 main states derived from k-means clustering of connectivity matrices from HC and ASD groups. (b) The corresponding connectome plots for HC and ASD.
Figures 2a and 2b show the t-distributed Stochastic Neighbor Embedding (t-SNE) visualizations of respective latent state sequences $\{\mathbf{Z}_{t}\}_{t=0}^{T}$ from models II and III of the ablation study. It shows that the latent state t-SNEs of model II are more concentrated towards the centre in comparison to the t-SNEs of model III, which is more uniform. This indicates that the VB elements (KLD and reparameterization) facilitate learning of a smooth latent embedding space to better accommodate subject-level spatiotemporal variability across unseen dynamic FC networks. Nevertheless, due to incorporation of the prior model via KLD loss (7), VB-based models generate latent embeddings that are highly accustomed to the causal structure of the dynamic FC data. This aggravates model overfitting and increases the chance of false classification, resulting in an inferior performance of model III in Table 2 compared to model II.
@@
In addition, we constructed higher-order dynamic FC by correlating the DSVB-learned embeddings $\mathbf{z}_{it}$ between pairs of nodes, and examined the dynamic connectivity states in ASD by applying dynamic FC clustering of these dynamic networks. The optimal number of clusters = 3 is determined using the Elbow method Bholowalia and Kumar (2014). Figure  3 shows the estimated connectivity matrices for three different states with the corresponding highest FC connections. By referring to Power et al. (2011), the corresponding brain functional system can be traced back. The ASD group shows stronger FC connections in visual, default mode, salience networks; whereas higher connectivity is observed in HC for sensory and auditory networks. The difference in FC patterns between ASD and HC is state-dependent. In ASD, visual ROIs show increased interconnections in state-1 and state-2 where default mode and salience are mostly activated in state-1. This is consistent with previous findings Holiga et al. (2019) that hyperconnectivity exists at prefrontal and parietal cortices; and hypoconnectivity restricted to sensory-motor regions for FC patterns in ASD as compared to HC. Figure 4 shows that the latent embedding matrices derived from DSVB are more visually discriminative compared to the raw FC matrices. The ASD embeddings highlight hyperconnectivity within visual networks (ROIs 143 to 173), particularly on the occipital lobe, and this aligns with earlier neuroscience discoveries Keehn et al. (2008); Clery et al. (2013); Matsuoka et al. (2020).
Figure 4: Comparison of the group means of raw FC with the latent embeddings derived from our DSVB framework over time for both HC and ASD groups.
@@ 5 Conclusion
Acknowledgement
The work of Chee-Ming Ting is supported by the Ministry of Higher Education, Malaysia under Fundamental Research Grant Scheme FRGS/1/2023/ICT02/MUSM/02/1. Support on compute resources was provided by the Advanced Performance Computing (ACP) Platform, Monash University Malaysia and Robust HPC, Malaysia. Sin-Yee Yap has been based at Monash’s Department of Data Science and AI for some of the duration of this work.

### v4

@@ Abstract
Recent applications of pattern recognition techniques on brain connectome classification using functional connectivity (FC) are shifting towards acknowledging the non-Euclidean topology and dynamic aspects of brain connectivity across time. In this paper, a deep spatiotemporal variational Bayes (DSVB) framework is proposed to learn time-varying topological structures in dynamic FC networks for identifying autism spectrum disorder (ASD) in human participants. The framework incorporates a spatial-aware recurrent neural network with an attention-based message passing scheme to capture rich spatiotemporal patterns across dynamic FC networks. To overcome model overfitting on limited training datasets, an adversarial training strategy is introduced to learn graph embedding models that generalize well to unseen brain networks. Evaluation on the ABIDE resting-state functional magnetic resonance imaging dataset shows that our proposed framework substantially outperforms state-of-the-art methods in identifying patients with ASD. Dynamic FC analyses with DSVB-learned embeddings reveal apparent group differences between ASD and healthy controls in brain network connectivity patterns and switching dynamics of brain states. The code is available at https://github.com/Monash-NeuroAI/Deep-Spatiotemporal-Variational-Bayes.
@@ 1 Introduction
In this paper, a deep spatiotemporal variational Bayes (DSVB) framework is proposed to learn the stochastic dynamic graph embeddings of FC networks, composed of sliding window segments of the resting-state fMRI time-series data. Figure 1 shows an overview of the proposed DSVB framework. Derived from general sequential variational Bayes (VB), our proposed framework provides a stochastic representation modelling solution of using dynamic FC networks for neuropsychiatric disorder diagnosis. Unlike existing GNN-based approaches that learn independent and identically distributed (i.i.d.) graph embeddings, we incorporate a spatial-aware recurrent neural network with attention-based message-passing scheme to model both the complex topological structure and temporal dependencies within a dynamic FC network sequence. Additionally, we introduce an adversarial training strategy to overcome the overfitting issue aggravated by limited training data. Our contributions are highlighted as follows:
@@
    An end-to-end probabilistic framework, integrating variational Bayes and recurrent GNNs, to learn the sequential stochastic graph latent embedding spaces in a purely unsupervised manner. This accommodates a wider range of spatiotemporal variability across unseen brain FC networks.
@@ 2.2 Spatiotemporal Variational Bayes
where $\mathbf{X}_{\leq t}$ and $\mathbf{X}_{<t}$ denote the partial sequences up to the $t^{\text{th}}$ and $({t-1})^{\text{th}}$ time samples, respectively. $\mathcal{D}^{\mathrm{KL}}$ denotes the (positive-valued) Kullback–Leibler divergence (KLD), which measures the statistical discrepancy of the encoder from the prior. The conditional probabilities of (3) encapsulate the underlying topological structure and temporal coherence of the dynamic brain networks. This sELBO (3) then forms the basis of our proposed DSVB framework.
@@ 2.5 Latent Embeddings for Graph Classification
modelled by multilayer FCNN. A softmax activation is then applied on the obtained logits $\mathbf{\hat{y}}\in\mathbb{R}^{2}$ to get the predictive probability scores for the final subject-level brain network classification. The predicted class label is thus the one with highest predictive score. The classification loss is taken to be the following multi-class cross entropy:
@@ 4 Results and Discussion
To evaluate the effectiveness of our proposed method, a nested-stratified 5-fold cross-validation was applied for each experiment. The performance of ASD identification is measured using five metrics: accuracy, recall, precision, F1-score, and area under the curve (AUC). Table 1 compares the average performances of the proposed DSVB-FCNN with the existing baseline classifiers based on both dynamic and static FC networks. It shows that our proposed DSVB-FCNN achieves substantial improvement over the baseline classifiers in each metric. The DSVB-FCNN also achieved low standard deviations in all metrics, demonstrating a strong consistency in its performances across the 5-fold cross-validation. In general, these results indicate that the framework facilitates robust brain disorder identification by generating graph latent embeddings that generalize well to unseen dynamic FC networks beyond the limited training dataset.
@@
Figure 3: The average connectivity patterns derived from DSVB-learned embeddings. (a) 3 main states derived from k-means clustering of 264 x 264 ROIs connectivity matrices from HC and ASD groups. (b) The corresponding connectome plots for HC and ASD.
Figures 2a and 2b show the t-distributed Stochastic Neighbor Embedding (t-SNE) visualizations of respective latent state sequences $\{\mathbf{Z}_{t}\}_{t=0}^{T}$ from models II and III of the ablation study. It shows that the latent state t-SNEs of model II are more concentrated towards the centre in comparison to the t-SNEs of model III, which is more uniform. This indicates that the VB elements (KLD and reparameterization) facilitate learning of a smooth latent embedding space to better accommodate subject-level spatiotemporal variability across unseen dynamic FC networks. Nevertheless, due to incorporation of the prior model via KLD loss (7), VB-based models generate latent embeddings that are highly accustomed to the spatiotemporal structure of the dynamic FC data. This aggravates model overfitting and increases the chance of false classification, resulting in an inferior performance of model III in Table 2 compared to model II.
@@
In addition, we constructed higher-order dynamic FC by correlating the DSVB-learned embeddings $\mathbf{z}_{it}$ between pairs of nodes, and examined the dynamic connectivity states in ASD by applying dynamic FC clustering of these dynamic networks. The optimal number of clusters = 3 is determined using the Elbow method Bholowalia and Kumar (2014). Figure  3 shows the estimated connectivity matrices for three different states with the corresponding highest FC connections. By referring to Power et al. (2011), the corresponding brain functional system can be traced back. The ASD group shows stronger FC connections in visual, default mode, salience networks; whereas higher connectivity is observed in HC for sensory and auditory networks. The difference in FC patterns between ASD and HC is state-dependent. In ASD, visual ROIs show increased interconnections in state-1 and state-2 where default mode and salience are mostly activated in state-1. This is consistent with Holiga et al. (2019), who also concluded that - compared to HC - ASD has hyperconnectivity in prefrontal and parietal cortices but hypoconnectivity in sensory-motor regions. Figure 4 shows that the latent embedding matrices derived from DSVB are more visually discriminative compared to the raw FC matrices. The ASD embeddings highlight hyperconnectivity within visual networks (ROIs 143 to 173), particularly on the occipital lobe, and this aligns with earlier neuroscience discoveries Keehn et al. (2008); Clery et al. (2013); Matsuoka et al. (2020).
Figure 4: Comparison of the group means of raw FC (264 x 264 ROIs) with the latent embeddings derived from our DSVB framework over time for both HC and ASD groups.
@@ 5 Conclusion
Acknowledgements
The work of Junn Yong Loo is supported by Monash University under the SIT Collaborative Research Seed Grants 2024 I-M010-SED-000242. The work of Chee-Ming Ting is supported by the Ministry of Higher Education, Malaysia under the Fundamental Research Grant Scheme FRGS/1/2023/ICT02/MUSM/02/1. Sin-Yee Yap has been based at Monash’s Department of Data Science and AI for some of the duration of this work.
Contribution Statement
Sin-Yee Yap, Junn Yong Loo (corresponding author) and Chee-Ming Ting contributed equally to this work.